### v1

# Uniform subelliptic estimates for degenerating Fokker-Planck equations

Abstract.
In this paper we expand upon recent work of Hernandez, Ranard and Riedel [5], [4], Galkowski and Zworski [3], and Li [7], by proving long time bounds for solutions to certain Fokker-Planck equations with subelliptic diffusion term. We consider the case where the jump operators $\ell_{j}$ in the Lindbladian are linear functions of $x$, and place an assumption which implies that the Hörmander condition holds for the resulting Fokker-Planck equation. By constructing a suitable parametrix for this equation we show that semiclassical derivative estimates established in [3] and [7] for elliptic diffusion also hold for subelliptic diffusion, with global bounds in $L^{p}$ for all $1\leq p\leq\infty$.
Key words and phrases:
Subelliptic, Fokker-Planck, Lindbladian
2020 Mathematics Subject Classification:
35Q84 (Primary), 35H20 (Secondary)
1. Introduction
Recent work by Hernandez, Ranard and Riedel [5], [4], Galkowski and Zworski [3], and Li [7], show that if a quantum system interacts suitably with an open environment, then the quantum evolution of an observable remains close to the classical evolution for times that greatly exceed the Ehrenfest time $t_{E}=\log(h^{-1})$. The observables are modeled as either semiclassical pseudodifferential operators or superpositions of Gaussian states, and the interaction with the environment by “jump” operators via the Lindblad master equation. The classical evolution is given by evolution of the symbol under the Fokker-Planck equation, which arises as the leading terms in the composition formula with the Lindbladian. We refer to the above works for a more detailed discussion of the physics, as well as numerical experiments that demonstrate the phenomenon of extended coherence.
The purpose of this note is to show that key results of these papers hold under relaxed assumptions concerning the interaction with the external environment. Roughly, in the above models the interaction involved diffusion that acted in both the space and momentum variables. The resulting Fokker-Planck equation then resembles the standard heat equation on phase space, with a small parameter in front of the diffusion that vanishes as $h\rightarrow 0$. In this work we establish the same key estimates for solutions to the Fokker-Planck equation assuming only diffusion in the momentum variables. In terms of the Lindblad jump operators, this means that the jump functions depend only on the spatial variable $x$. We assume that the diffusion spans the momentum directions, which yields that the Fokker-Planck operator is a subelliptic heat equation of Hörmander sum-of-squares type [6].
The focus of this paper is establishing the estimates of Theorem 1.1 for the Fokker-Planck equation, which we do by producing a suitable parametrix to any given order. The paremetrix is constructed via a non-isotropic heat kernel iteration, analogous to the isotropic iteration in [7]. This allows us to establish bounds in $L^{p}$ for all $1\leq p\leq\infty$. Estimates like Theorem 1.1 for $p=2$ and $p=1$ are key to the conclusions of [3] and [7], and in the final section we show how certain of their results also hold in the subelliptic case.
Similar to the above papers, we consider as given a Schrödinger type Hamiltonian $p(x,\xi)=\frac{1}{2}|\xi|^{2}+V(x)$, which represents the closed quantum system. We will assume $V(x)$ is smooth and real-valued function on $\mathbb{R}^{n}$, and also assume that
(1.1) $|\partial_{x}^{\alpha}V(x)|\leq\begin{cases}C\,\langle x\rangle^{2-|\alpha|},&|\alpha|\leq 2,\\ C_{\alpha},&|\alpha|\geq 2.\end{cases}$
We use $H_{p}$ to denote the corresponding Hamiltonian vector field
$H_{p}=\xi\cdot\partial_{x}-\nabla V(x)\cdot\partial_{\xi}.$
As in [5] we consider a family of linear jump operators, but assume that $\{\ell_{j}(x)\}_{j=1}^{J}$ depend only on the space variable $x$,
$\ell_{j}(x)=\sum_{i=1}^{n}c_{ji}\,x_{i},\quad 1\leq j\leq J,\quad c_{ji}\in\mathbb{C}.$
If $L_{j}$ denotes multiplication by $\ell_{j}(x)$, then $L_{j},L_{j}^{*}]=0=\{\ell_{j},\bar{\ell}_{j}\}$, and the Lindbladian with coupling constant $\gamma$ simplifies to
$\mathcal{L}A=\frac{i}{h}[P,A]+\frac{\gamma}{2h}\sum_{j=1}^{J}L_{j}[A,L_{j}^{*}]+[L_{j},A]L_{j}^{*}.$
This is a model for the evolution of an observable $A$ developing under the Schrödinger system interacting stochastically with a larger environment through the collection of jump operators $L_{j}$.
The corresponding Fokker-Planck operator, which gives the leading term in $\mathcal{L}A$ when $P=p^{w}(x,hD)$ and $A=a^{w}(x,hD)$, is
(1.2) $Q=H_{p}+\gamma\sum_{j=1}^{J}\mathop{\rm Im}\nolimits(\bar{\ell}_{j}\nabla_{x}\ell_{j})\cdot\partial_{\xi}+\frac{\gamma h}{2}\sum_{j=1}^{J}(\nabla_{x}\ell_{j}\cdot\partial_{\xi})(\nabla_{x}\bar{\ell}_{j}\cdot\partial_{\xi}).$
We assume that the second order terms are non-degenerate in $\partial_{\xi}$, which implies that there is a real, non-singular $n\times n$ matrix $B$,
$\sum_{j=1}^{J}\mathop{\rm Re}\nolimits\Bigl((\nabla_{x}\ell_{j})\otimes(\nabla_{x}\bar{\ell}_{j})\Bigr)=BB^{T}.$
We may then write
(1.3) $\sum_{j=1}^{J}(\nabla_{x}\ell_{j}\cdot\partial_{\xi})(\nabla_{x}\bar{\ell}_{j}\cdot\partial_{\xi})=\sum_{j=1}^{n}X_{j}^{2},\quad X_{j}=\sum_{i=1}^{n}B_{ij}\partial_{\xi_{i}}.$
The main result of this paper concerns uniform regularity estimates for the solution $u=e^{tQ}u_{0}$ to the initial value problem for the Fokker-Planck equation; for $t>0$:
$\partial_{t}u=Qu,\qquad u(0,x,\xi)=u_{0}(x,\xi).$
We assume that both $\gamma,h\in(0,1]$, and let $\epsilon=\sqrt{\gamma h/2}$. If we treat $\epsilon$ and $\gamma$ as the independent variables then the estimates involve $\epsilon$, but the constants in the estimates are uniform over $\gamma,\epsilon\in(0,1]$. Li [7] considered the case that $1\leq\gamma\leq h^{-1}$, but the presence of the linear term in $\partial_{\xi}$ in (1.2), i.e.  the first sum in (1.2), leads us to require bounds on $\gamma$. For simplicity we assume $\gamma\leq 1$. If $\ell_{j}$ is real valued, however, the estimates of Theorem 1.1 hold for $\gamma\leq h^{-1}$ since in that case $\epsilon\leq 1$ and the term involving just $\gamma$ in (1.2) goes away.
Since the second order terms in $Q$ are constant coefficient, the derivative bounds on $V$ and $\ell_{j}(x)$ show that $Q,(\epsilon\partial_{x})^{\alpha}(\epsilon\partial_{\xi})^{\beta}]$ is a differential operator in $\epsilon\partial_{x}$ and $\epsilon\partial_{\xi}$ of order at most $|\alpha|+|\beta|$, with smooth bounded coefficients. The operator $e^{tQ}$ is a contraction on $L^{p}$ for all $1\leq p\leq\infty$ (see Section 6). Thus, by the Duhamel formula and Gronwall’s inequality we conclude that, for all $N$ and all $\epsilon\in[0,1]$,
(1.4) $\sum_{|\alpha|+|\beta|\leq N}\|(\epsilon\partial_{x})^{\alpha}(\epsilon\partial_{\xi})^{\beta}u(t,\cdot)\|_{L^{p}(\mathbb{R}^{n})}\\ \leq C_{N}e^{ct}\sum_{|\alpha|+|\beta|\leq N}\|(\epsilon\partial_{x})^{\alpha}(\epsilon\partial_{\xi})^{\beta}u_{0}\|_{L^{p}(\mathbb{R}^{n})}.$
Our result, analogous to ones in [3] and [7], is that these bounds hold with a constant independent of $t$, i.e.  $c=0$. The following theorem counts the order of derivatives in $x$ and $\xi$ as they are counted in the subelliptic calculus, but combining (1.4) for $t\leq 1$ with (1.6) for $T=1$ shows that (1.4) holds as written with $c=0$, for some $C_{N}$ independent of $t$.
Theorem 1.1.
Assume that $p=\frac{1}{2}|\xi|^{2}+V(x)$ with real $V$ satisfying (1.1), and $Q$ is of the form (1.2), where (1.3) holds with non-singular $B$. Then for all $N\in\mathbb{Z}_{+}$ there is $C_{N}$ such that, with $u=e^{tQ}u_{0}$ and $\epsilon=\sqrt{\gamma h/2}$, for all $1\leq p\leq\infty$, and all $\epsilon,\gamma\in(0,1]$,
(1.5) $\sup_{0\leq t\leq\infty}\,\sum_{|\alpha|+3|\beta|\leq 3N}\|(\epsilon\partial_{\xi})^{\alpha}(\epsilon\partial_{x})^{\beta}u(t,\cdot)\|_{L^{p}(\mathbb{R}^{n})}\\ \leq C_{N}\sum_{|\alpha|+3|\beta|\leq 3N}\|(\epsilon\partial_{\xi})^{\alpha}(\epsilon\partial_{x})^{\beta}u_{0}\|_{L^{p}(\mathbb{R}^{n})}.$
Additionally, for all $N\in\mathbb{Z}_{+}$ and $T>0$ there is $C_{N,T}$ so that
(1.6) $\sup_{T\leq t\leq\infty}\,\sum_{|\alpha|+3|\beta|\leq N}\|(\epsilon\partial_{\xi})^{\alpha}(\epsilon\partial_{x})^{\beta}u(t,\cdot)\|_{L^{p}(\mathbb{R}^{n})}\leq C_{N,T}\|u_{0}\|_{L^{p}(\mathbb{R}^{n})},$
where $C_{N,T}=\mathcal{O}(T^{-\frac{1}{2}N})$ as $T\rightarrow 0^{+}$.
Estimate (1.6) captures the smoothing effect of $e^{tQ}$ for $t>0$. Estimate (1.5) is the result of the parametrix mapping properties, where the counting of $\alpha$ and $\beta$ reflects the relative order of $\partial_{x}$ versus $\partial_{\xi}$ in the nonisotropic subelliptic calculus associated to $Q$.
Our proof of Theorem 1.1 is a modification of our earlier work [9]. Unlike in that work, the Fokker-Planck equation has a distinguished time variable, which we exploit to handle the non-compact setting. Estimates are obtained through a parametrix construction, which we outline here in the case that $B$ is the identity matrix.
We start by pulling the Fokker-Planck equation back by the flow along the first-order terms in (1.2). The leading order terms in the pullback of $\partial_{t}-Q$ are then $\partial_{t}-(\partial_{\xi}+t\partial_{x})^{2}$, in the subelliptic notion of order. It is convenient to make a linear change of variables to put this in the form
$\partial_{t}-{\tfrac{1}{2}}\xi\cdot\partial_{x}-(\partial_{\xi}+{\tfrac{1}{2}}t\partial_{x})^{2},$
which is a Hörmander type operator involving left-invariant vector fields on a step 2 nilpotent Lie group. We will see that the fundamental solution on this Lie group can be written explicitly in terms of the Gaussian heat kernel.
This produces the leading term in a left/right inverse for $\partial_{t}-Q$, and we then use heat kernel iteration steps to produce a left or right inverse to any desired order. In the iteration step composition of kernels is expressed as convolution in the group product.
Care needs to be taken in that the pullback of $\partial_{\xi}$ and $\partial_{x}$ by the flow gives a frame whose second derivatives in $t$ increase unboundedly with $\xi$ due to the $\xi\cdot\partial_{x}$ term in the flow. To obtain global estimates we therefore can use only the first few derivatives in $t$, which we accomplish by using the heat kernel form of the fundamental solution rather than the pseudodifferential representation used in [9].
2. The model operator and the case $\epsilon=1$
We work with the following frame of vector fields on $\mathbb{R}^{2n+1}$, in the variable $y=(y_{0},y^{\prime},y^{\prime\prime})\in\mathbb{R}\times\mathbb{R}^{n}\times\mathbb{R}^{n}$,
  • $\bullet$
    $Y_{0}=\partial_{0}-{\frac{1}{2}}\sum_{j=1}^{n}y_{j}\partial_{j+n}$
  • $\bullet$
    $Y_{j}=\partial_{j}+{\frac{1}{2}}y_{0}\partial_{j+n}\;$ for  $1\leq j\leq n\rule${0.0pt}{17.0pt},
  • $\bullet$
    $Y_{j}=\partial_{j}\;$ for $j\geq n+1\rule${0.0pt}{17.0pt}.
These satisfy the commutation relations
$Y_{0},Y_{j}]=Y_{j+n}\quad\text{if}\quad 1\leq j\leq n,$
with all other commutators equal to 0. The $Y_{j}$ form a step-2 nilpotent Lie algebra, and are the left-invariant frame of vector fields associated to the Lie group structure on $\mathbb{R}^{2n+1}$ with product
$z\cdot y=(y_{0}+z_{0},y^{\prime}+z^{\prime}\!,y^{\prime\prime}+z^{\prime\prime}+{\tfrac{1}{2}}z_{0}y^{\prime}-{\tfrac{1}{2}}y_{0}z^{\prime}).$
Observe that $z^{-1}=-z$. We will use the shorthand notation
$Y^{\prime}=(Y_{1},\ldots,Y_{n}),\qquad Y^{\prime\prime}=(Y_{n+1},\ldots,Y_{2n}).$
There is an associated dilation structure, given by
$\delta_{r}(y)=\bigl(r^{2}y_{0},ry^{\prime},r^{3}y^{\prime\prime}\bigr),$
under which $_{0}, YY^{\prime}$, and $Y^{\prime\prime}$ are respectively of order $2$, 11, and $3$, where we say an operator $L$ is of order $j$ if
$L(f\circ\delta_{r})=r^{j}(Lf)\circ\delta_{r}\quad\text{for}\;\,r>0.$
For a multi-index $\alpha$, we define its order to be
$\operatorname{order}(\alpha)=2\alpha_{0}+\alpha_{1}+\cdots+\alpha_{d}+3\alpha_{d+1}+\cdots+3\alpha_{2d}=2\alpha_{0}+|\alpha^{\prime}|+3|\alpha^{\prime\prime}|.$
Consider $\alpha\in\mathbb{Z}\times\mathbb{Z}_{+}^{2n}$, $\beta\in\mathbb{Z}_{+}^{2n+1}$. That is, we will allow $\alpha_{0}<0$, but require $\alpha^{\prime},\alpha^{\prime\prime}\geq 0$ and $\beta\geq 0$. Then we have
$\operatorname{order}(y^{\alpha}Y^{\beta})=\operatorname{order}(y^{\alpha}\partial_{y}^{\beta})=\operatorname{order}(\beta)-\operatorname{order}(\alpha).$
Observe that $y^{\alpha}Y^{\beta}$ can be written as a linear combination of terms of the form $y^{\theta}\partial_{y}^{\gamma}$ of the same order, and vice-versa.
The operator
$P_{0}(y,\partial_{y})=\Bigl(Y_{0}-\sum_{j=1}^{n}Y_{j}^{2}\Bigr)$
is subelliptic since it satisfies the Hörmander condition, see [6] and [8]. By a result of Folland [2], it has a unique fundamental solution that is homogeneous of degree 2-2. In our case there is an elementary formula for $K_{0}(y)$,
$K_{0}(y)=G_{y_{0}}(y^{\prime})G_{\frac{1}{12}y_{0}^{3}}(y^{\prime\prime}),$
where $G=e^{t\Delta}$ is the standard heat kernel on $\mathbb{R}^{n}$:
$G_{t}(w)={\mathds{1}}_{t>0}\,(4\pi t)^{-\frac{n}{2}}\;e^{-\frac{|w|^{2}}{4t}}.$
The solution to the initial value problem $P_{0}(y,\partial_{y})u=0$ for $y_{0}>0$ with $u|_{y_{0}=0}=f$ is given by
$u(y)=\int K_{0}(y_{0},y^{\prime}-z^{\prime},y^{\prime\prime}-z^{\prime\prime}+{\tfrac{1}{2}}y_{0}z^{\prime})f(z^{\prime},z^{\prime\prime})\,dz^{\prime}\,dz^{\prime\prime}.$
Definition 2.1.
For $\gamma\in\mathbb{Z}\times\mathbb{Z}_{+}^{2n}$, we define $K_{\gamma}(y)=y^{\gamma}K_{0}(y)$. We say that $K(y)$ is a kernel of order $m$, respectively of order $\leq m$, if $K$ is a finite linear combination $\sum c_{\gamma}K_{\gamma}$ with $\operatorname{order}(\gamma)=-m$, respectively $\operatorname{order}(\gamma)\geq-m$.
Observe that for each $\alpha\in\mathbb{Z}_{+}^{2n+1}$ one can write
(2.1) $\partial_{y}^{\beta}K_{0}(y)=\sum c_{\beta,\gamma}\,K_{\gamma}(y),$
where $\gamma^{\prime},\gamma^{\prime\prime}\geq 0$, $\operatorname{order}(\gamma)=-\operatorname{order}(\beta)$, and $0\geq\gamma_{0}\geq-3|\beta|$. That is, $\partial_{y}^{\beta}K_{0}$ is a kernel of order equal to $\operatorname{order}(\beta)$. The same holds for $Y^{\beta}K_{0}$.
Lemma 2.2.
If $\gamma\in\mathbb{Z}\times\mathbb{Z}_{+}^{2n}$ then one can write
$K_{\gamma}(y)=\sum_{j,\theta}b_{j,\theta}\,y_{0}^{j}\,\partial_{y}^{\theta}K_{0}(y)=\sum_{j,\theta}c_{j,\theta}\,y_{0}^{j}\,Y^{\theta}K_{0}(y),$
where $\theta_{0}=0$, $|\theta|\leq|\gamma|$ and $\operatorname{order}(y_{0}^{j}\partial_{y}^{\theta})=\operatorname{order}(y_{0}^{j}Y^{\theta})=\operatorname{order}(y^{\gamma})$.
Proof.
The result follows by observing that
$y^{\prime\prime}K_{0}=-\frac{1}{6}\,y_{0}^{3}\,Y^{\prime\prime}K_{0},\qquad y^{\prime}K_{0}=-2y_{0}\partial_{y^{\prime}}K_{0}=(y_{0}^{2}Y^{\prime\prime}-2y_{0}Y^{\prime})K_{0},$
and using the fact that commutation between monomials in $y$ and monomials in $Y$ preserves the order. ∎
Lemma 2.3.
Assume $f(s)\in C([0,\infty))$, and let $F(z_{0})=\displaystyle\int_{0}^{z_{0}}f(s)\,ds$. Then
$\int K_{0}(z^{-1}y)f(z_{0})K_{0}(z)\,dz=F(y_{0})K_{0}(y).$
Proof.
One can be verify by calculation that, for each $y_{0},z_{0}$,
$\int K_{0}(z^{-1}y)K_{0}(z)\,dz^{\prime}\,dz^{\prime\prime}=K_{0}(y).$
The statement also follows from the fact that applying $Y_{0}-\sum_{j=1}^{n}Y_{j}^{2}$ yields $f(y_{0})K_{0}(y)$ on both sides, and both sides vanish for $y_{0}<0$. ∎
Corollary 2.4.
Suppose that $\operatorname{order}(\gamma)\geq-1$, and $f_{\gamma}\in C([0,\infty))$. Then one can write
$\int K_{0}(z^{-1}y)f_{\gamma}(z_{0})K_{\gamma}(z)\,dz=\sum_{\theta}f_{\theta}(y_{0})K_{\theta}(y),$
where the sum is finite, $\operatorname{order}(\theta)=\operatorname{order}(\gamma)+2$, and $f_{\theta}\in C([0,\infty))$.
Proof.
We use Lemma 2.2 to write the integral as a finite sum with $\theta_{0}=0$,
$\sum_{j,\theta}b_{j,\theta}\int K_{0}(z^{-1}y)z_{0}^{j}f_{\gamma}(z_{0})\partial_{z}^{\theta}K_{0}(z)\,dz$
where $j\geq 0$ since $\operatorname{order}(z_{0}^{j}\partial_{z}^{\theta})=-\operatorname{order}(\gamma)\leq 1$. We integrate by parts in $(z^{\prime},z^{\prime\prime})$, and use that
$\partial_{z^{\prime\prime}}K_{0}(z^{-1}y)=\partial_{y^{\prime\prime}}K_{0}(z^{-1}y),\quad-\partial_{z^{\prime}}K_{0}(z^{-1}y)=(\partial_{y^{\prime}}-y_{0}\partial_{y^{\prime\prime}})K_{0}(z^{-1}y).$
The integral is thus a finite sum
$\sum d_{i,j,\theta}\,y_{0}^{i}\partial_{y}^{\theta}\int K_{0}(z^{-1}y)z_{0}^{j}f_{\gamma}(z_{0})K_{0}(z)\,dz,$
where the sum is now over $\operatorname{order}(y_{0}^{i}\partial_{y}^{\theta})=2j-\operatorname{order}(\gamma)$ and $\theta_{0}=0$. We apply Lemma 2.3, and write
$\int_{0}^{y_{0}}s^{j}f_{\gamma}(s)\,ds=y_{0}^{j+1}\int_{0}^{1}s^{j}f_{\gamma}(st)\,ds.$
The proof is concluded by applying (2.1). ∎
We will see that in flow coordinates the Fokker-Planck equation is given by a differential operator in $y$, with coefficients that are smooth in $(y^{\prime},y^{\prime\prime})$, but with only finitely many bounded derivatives in $y_{0}$. After an appropriate Taylor expansion in $(y^{\prime},y^{\prime\prime})$, the Fokker-Planck operator is then approximated to a prescribed order by a sum
$P_{0}(y,\partial_{y})+\sum_{j=1}^{N}P_{j}(y,\partial_{y}),$
where $P_{j}(y,\partial_{y})$ is a differential operator in $(\partial_{y^{\prime}},\partial_{y^{\prime\prime}})$ of the form
(2.2) $P_{j}(y,\partial_{y})=\sum_{\alpha,\beta}f_{\alpha,\beta}(y_{0})y^{\alpha}\partial_{y}^{\beta}\;:\;\beta_{0}=0,\;\alpha\geq 0,\;\operatorname{order}(y^{\alpha}\partial_{y}^{\beta})=2-j,$
with $f_{\alpha,\beta}\in C^{0}(\mathbb{R})$. The functions $f_{\alpha,\beta}(y_{0})$ are in fact smooth, but we have uniform control only on $\sup_{|y_{0}|<T}|f_{\alpha,\beta}(y_{0})|$ for each $T$, and not on derivatives in $y_{0}$. In our application $|\beta|\leq 2$, but that is not needed for the following.
Theorem 2.5.
Assume given operators $P_{j}(y,\partial_{y})$ of the form (2.2) for $j\geq 1$. Then there are kernels $K_{m}$ and $R_{N}$ of the form
$K_{m}(y)=\sum_{\operatorname{order}(\gamma)=m}b_{\gamma}(y_{0})K_{\gamma}(y),\quad R_{N}(y)=\sum_{\operatorname{order}(\gamma)\geq N-1}r_{\gamma}(y_{0})K_{\gamma}(y),$
with $b_{\gamma},r_{\gamma}\in C(\mathbb{R})$, so that for each NN
$\Bigl(P_{0}(y,\partial_{y})+\sum_{j=1}^{N}P_{j}(y,\partial_{y})\Bigr)\Bigl(K_{0}(y)+\sum_{m=1}^{N}K_{m}(y)\Bigr)=\delta(y)+R_{N}(y).$
Proof.
We recursively solve $P_{0}K_{m}=-\sum_{j=1}^{m}P_{j}K_{m-j}$ by letting
$K_{m}(y)=-\int K_{0}(z^{-1}y)\Bigl(\sum_{j=1}^{m}P_{j}K_{m-j}\Bigr)(z)\,dz.$
By Corollary 2.4 this leads to $K_{m}$ of the stated form, and we then have
$R_{N}=\sum_{\begin{subarray}{c}j,m\leq N\\ j+m\geq N+1\end{subarray}}P_{j}K_{m}$
which yields a sum of terms of the stated form. ∎
The recursion shows that, for any $T<\infty$, if each $f_{\alpha,\beta}\in C([0,T])$ then so are $b_{\gamma}$ and $r_{\gamma}$, and one can bound $\sup_{y_{0}\leq T}|b_{\gamma}(y_{0})|$ and $\sup_{y_{0}\leq T}|r_{\gamma}(y_{0})|$ from upper bounds on $\sup_{y_{0}\leq T}|f_{\alpha,\beta}(y_{0})|$ and $T$.
3. The Fokker-Planck equation
In this section we represent the Fokker-Planck operator as a sum of squares, and use the flow along $X_{0}$ to map the fundamental solution of Section 2 to the $(t,x,\xi)$ variables. To keep notation concise, we will make use of $w=(t,x,\xi)$ and $v=(s,z,\zeta)$ to denote variables on $\mathbb{R}^{2n+1}$. In relating the Fokker-Planck operator to the Lie group frame $Y_{j}$, $\partial_{x}$ will be comparable to $Y^{\prime\prime}=\{Y_{j}\}_{j=n+1}^{2n}$, and $\partial_{\xi}$ to $Y^{\prime}=\{Y_{j}\}_{j=1}^{n}$. Thus it is natural to write $w=(w_{0},w^{\prime\prime},w^{\prime})$, or $t=w_{0}$, $x=w^{\prime\prime}$, and $\xi=w^{\prime}$.
Consider the following frame of vector fields on $\mathbb{R}^{2n+1}$, where the index $j$ runs over $1\leq j\leq n$, and the $X_{j}=\sum_{i=1}^{n}B_{ji}\cdot\partial_{\xi_{i}}$ are as in (1.3),
[…]
The choice of $X_{j+n}$ is such that $X_{0},X_{j}]=X_{j+n}$.
Let $\Phi_{t}:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$ denote the flow map along $X_{0}$ for time $t$. Since $X_{0}$ is divergence free, $\Phi_{t}$ preserves the volume form $dx\wedge d\xi$ for each $t$. From the global $C^{\infty}$ bounds $|\partial_{x,\xi}^{\alpha}X_{0}|\leq C_{\alpha}$ for all $|\alpha|\geq 1$, it follows that the map $\Phi_{t}$ is globally bilipschitz on $\mathbb{R}^{2n}$, with Lipschitz constants bounded uniformly over $t\in[-T,T]$ for each $T>0$. Additionally, for each $|\alpha|\geq 1$ there are uniform bounds
(3.1) $\sup_{x,\xi}|\partial_{x,\xi}^{\alpha}\Phi_{t}(x,\xi)|\leq C_{\alpha}(T)\quad\text{for all}\;\;t\in[-T,T].$
Let $\{Z_{j}\}_{j=1}^{2n}=(\Phi_{-t})_{*}\{X_{j}\}_{j=1}^{2n}$ be the the pullback of $\{X_{j}\}_{j=1}^{2n}$ under $\Phi_{t}$. Then $\{Z_{j}\}_{j=1}^{2n}$ is a smooth frame on $\mathbb{R}^{2n}$, and if we set
$W(t,x,\xi)=B^{T}\cdot\Bigl[(\nabla^{2}V)(\Phi_{t}(x,\xi))+\gamma\sum_{j=1}^{J}\mathop{\rm Im}\nolimits\bigl(\nabla\bar{\ell}_{j}\otimes\nabla\ell_{j}\bigr)\Bigr]\cdot(B^{T})^{-1},$
then for $1\leq j\leq n$,
$\partial_{t}Z_{j}=Z_{j+n},\qquad\partial_{t}Z_{j+n}=-\sum_{k=1}^{n}W_{jk}(t,x,\xi)Z_{k}.$
Although $|\partial_{t}W(t,x,\xi)|\sim|\xi\cdot(\nabla_{x}^{3}V)(\Phi_{t}(x,\xi))|$ is not globally bounded unless $V$ is quadratic, we know from the bounds (1.1) and (3.1) that
$|\partial_{x}^{\beta}\partial_{\xi}^{\alpha}W(t,x,\xi)|\leq C_{\alpha,\beta}(T)\quad\text{for all}\;\;t\in[-T,T].$
From the above equation we deduce that for all $\alpha,\beta$, and $1\leq j\leq n$,
$\sup_{k\leq 2}|\partial_{t}^{k}\partial_{x}^{\beta}\partial_{\xi}^{\alpha}Z_{j}|+\sup_{k\leq 1}|\partial_{t}^{k}\partial_{x}^{\beta}\partial_{\xi}^{\alpha}Z_{j+n}|\leq C_{\alpha,\beta}(T)\quad\text{for}\;\;t\in[-T,T].$
Definition 3.1.
We define $C^{j}_{t}C^{\infty}_{x,\xi}$ to be the space of functions $f$ on $\mathbb{R}^{2n+1}$ such that, for every $T$ and $\alpha\in\mathbb{Z}_{+}^{2n}$, there are bounds
$\sum_{i=0}^{j}\,\sup_{|t|\leq T}\,\sup_{x,\xi}|\partial_{t}^{i}\partial_{x,\xi}^{\alpha}f(t,x,\xi)|\leq C_{\alpha}(T).$
Similar notation applies when $(x,\xi)$ is replaced by a variable in $\mathbb{R}^{m}$.
Consequently, for $1\leq j\leq n$,
$Z_{j}$ $=(B^{T}\partial_{\xi})_{j}+t(B^{T}\partial_{x})_{j}+A_{2}(t,x,\xi)\cdot(\partial_{\xi},\partial_{x}),$
$Z_{j+n}$ $=(B^{T}\partial_{x})_{j}+A_{1}(t,x,\xi)\cdot(\partial_{\xi},\partial_{x}),\rule${0.0pt}{15.0pt}
where $A_{j}$ is a $(2n)^{2}$ matrix of functions in $C^{j}_{t}C^{\infty}_{x,\xi}$, with
$A_{j}(t,x,\xi)=\mathcal{O}(|t|^{j})\quad\text{as}\quad t\rightarrow 0.$
[…]
$t=y_{0},\quad\xi=\tilde{\xi}+By^{\prime},\quad x=\tilde{x}+By^{\prime\prime}+\tfrac{1}{2}y_{0}By^{\prime}.$
Then
$\partial_{y_{0}}-{\tfrac{1}{2}}y^{\prime}\partial_{y^{\prime\prime}}$ $=\partial_{t}$
$\partial_{y^{\prime}}+{\tfrac{1}{2}}y_{0}\partial_{y^{\prime\prime}}$ $=B^{T}\partial_{\xi}+tB^{T}\partial_{x}\rule${0.0pt}{12.0pt}
$\partial_{y^{\prime\prime}}$ $=B^{T}\partial_{x}\rule${0.0pt}{12.0pt}
[…]
Lemma 3.2.
For $w=(t,x,\xi)\in\mathbb{R}^{2n}$ define
$\exp_{w}(y)=\bigl(t+y_{0},\Phi_{y_{0}}(x+By^{\prime\prime}+\tfrac{1}{2}y_{0}By^{\prime},\xi+By^{\prime})\bigr).$
Then with $Y_{j}$ the left invariant vector fields defined above, the pullback of $X_{j}$ to the $y$ coordinates satisfies
$X_{0}=Y_{0},\quad X^{\prime}=Y^{\prime}+A_{2}(y)\cdot(\partial_{y^{\prime}},\partial_{y^{\prime\prime}}),\quad X^{\prime\prime}=Y^{\prime\prime}+A_{1}(y)\cdot(\partial_{y^{\prime}},\partial_{y^{\prime\prime}}),$
where the $A_{j}(y)$ are $n\times 2n$ matrix functions with coefficients in $C^{j}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$, and $\partial_{y_{0}}^{i}A_{j}(0,y^{\prime},y^{\prime\prime})=0$ for $0\leq i<j$.
Remark $3.$
$A_{j}(y)$ depends on $(x,\xi)$, although not $t$. Differentiating $\exp_{w}(y)$ in $(x,\xi)$ is equivalent to differentiating in $(y^{\prime},y^{\prime\prime})$, so we have the same bounds for derivatives of $A_{j}$ in $(x,\xi)$.
Definition 3.4.
Let $\Theta_{w}(v)$ be the inverse to $\exp_{w}(y)$:
$y=\Theta_{w}(v)\quad\Leftrightarrow\quad\exp_{w}(y)=v.$
With $w=(t,x,\xi)$ and $v=(s,z,\zeta)$, this equates to
$y_{0}=s-t,\quad(By^{\prime},By^{\prime\prime}+\tfrac{1}{2}(s-t)By^{\prime})=\Phi_{t-s}(z,\zeta)-(x,\xi).$
Thus if we let $\Phi_{s-t}(z,\zeta)=(z_{t-s},\zeta_{t-s})$, then
$\Theta_{w}(v)=\bigl(s-t,B^{-1}(\xi-\zeta_{t-s}),B^{-1}(x-z_{t-s}+{\tfrac{1}{2}}(t-s)(\xi-\zeta_{t-s}))\bigr)$
Lemma 3.2 states that, with $X_{j}$ acting in $v$ for fixed $w$,
$X^{\prime}\bigl(f(\Theta_{w}(v)\bigr)$ $=\bigl(Y^{\prime}f+A_{2}(y)\cdot(\partial_{y^{\prime}}f,\partial_{y^{\prime\prime}}f)\bigr)(\Theta_{w}(v)),$
$X^{\prime\prime}\bigl(f(\Theta_{w}(\cdot)\bigr)$ $=\bigl(Y^{\prime\prime}f+A_{1}(y)\cdot(\partial_{y^{\prime}}f,\partial_{y^{\prime\prime}}f)\bigr)(\Theta_{w}(\cdot)).$
With $X_{j}$ acting in $w$ for fixed $v$, we have the more direct connection
(3.2) $X_{j}\bigl(f(-\Theta_{w}(v)\bigr)=(Y_{j}f)(-\Theta_{w}(v)),\quad j\geq 1.$
Since $A_{2}(0,y^{\prime},y^{\prime\prime})=\partial_{y_{0}}A_{2}(0,y^{\prime},y^{\prime\prime})=0$, we can write $A_{2}(y)=y_{0}^{2}A_{0}(y)$ and $A_{1}(y)=y_{0}A_{0}(y)$, in each case with $A_{0}\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$.
In exponential coordinates at a given point $w$, for $1\leq i\leq 2n$ we take a Taylor expansion to order $N$ in $(y^{\prime},y^{\prime\prime})$ about $y^{\prime}=y^{\prime\prime}=0$ to write
(3.3) $X_{i}=Y_{i}+\sum_{k=1}^{2n}\biggl(\sum_{|\alpha|\leq N+2}c_{i,\alpha,k}(y_{0})\,y^{\alpha}\partial_{k}+\sum_{|\alpha|=N+3}r_{i,\alpha,k}(y)\,y^{\alpha}\partial_{k}\biggr)$
Here $\alpha_{0}=2$ if $1\leq i\leq n$, and $\alpha_{0}=1$ if $n+1\leq i\leq 2n$, so $\operatorname{order}(y^{\alpha}\partial_{k})\leq\operatorname{order}(Y_{i})-2$ in the sums. The functions $c_{i,\alpha,k}(y_{0})\in C^{0}_{y_{0}}C^{\infty}_{w}$ are continuous in $y_{0}$. They depend smoothly on $w$, but satisfy uniform bounds on [T,T][-T,T] for each $T$. Also, $r_{j,\alpha,k}(y)\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$. Note that $\operatorname{order}(y^{\alpha}\partial_{k})\leq-N$ if $|\alpha|\geq N+3$.
Expressing $\partial_{k}$ for $k\geq 1$ in terms of $Y_{k}$ with $k\geq 1$ yields an expansion with similar $c_{i,\alpha,k}(y_{0})$ and $r_{j,\alpha,k}(y)$ and $\operatorname{order}(y^{\alpha}Y_{k})\leq\operatorname{order}(Y_{i})-2$,
$X_{i}=Y_{i}+\sum_{k=1}^{2n}\biggl(\sum_{|\alpha|\leq N+2}c_{i,\alpha,k}(y_{0})\,y^{\alpha}Y_{k}+\sum_{|\alpha|=N+3}r_{i,\alpha,k}(y)\,y^{\alpha}\partial_{k}\biggr).$
This relation is symmetric, so we have the reverse expansion with different $c_{i,\alpha,k}\in C^{0}C^{\infty}_{w}$ and $r_{i,\alpha,k}\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$, and $\operatorname{order}(y^{\alpha}X_{k})\leq\operatorname{order}(Y_{i})-2$,
(3.4) $Y_{i}=X_{i}+\sum_{k=1}^{2n}\biggl(\sum_{|\alpha|\leq N+2}c_{i,\alpha,k}(y_{0})\,y^{\alpha}X_{k}+\sum_{|\alpha|=N+3}r_{i,\alpha,k}(y)\,y^{\alpha}\partial_{k}\biggr).$
Let $P_{j}$ be the terms of order 2j2-j in the expansion of $X_{0}-\sum_{i=1}^{n}X_{i}^{2}$ using (3.3), where $P_{0}=Y_{0}-\sum_{j=1}^{n}Y_{j}^{2}$. Following Theorem 2.5 we can iteratively construct $K_{m}(w,y)$ of the form
$K_{m}(w,y)=\sum_{\operatorname{order}(\gamma)=m}b_{\gamma}(y_{0},w)\,K_{\gamma}(y),$
so that, with $X_{j}$ written as a differential operator in $y$,
$\Bigl(X_{0}-\sum_{i=1}^{n}X_{i}^{2}\Bigr)\Bigl(K_{0}(y)+\sum_{m=1}^{N}K_{m}(w,y)\Bigr)=\delta(y)+R_{N}(w,y),$
where $\operatorname{order}(R_{N})\leq 1-N$, and $b_{\gamma}\in C^{0}_{y_{0}}C^{\infty}_{w}$ and $r_{\gamma}(y_{0},y,w)\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$. Both are independent of $t$.
We first use this to produce a right inverse for $X_{0}-\sum_{i=1}^{n}X_{i}^{2}$ modulo an operator of any given negative order.
Given $N$, let $K(v,y)=\sum_{m=0}^{N}K_{m}(v,y)$, and define
$T_{K}f(w)=c_{B}\int K(v,\Theta_{v}(w))\,f(v)\,\chi(\Theta_{v}(w))\,dv,$
where $c_{B}^{-1}=|D_{y}\exp_{v}(y)|_{y=0}=\det(B)^{2}$ and $\chi(y)\in C_{c}^{\infty}(\mathbb{R}^{2n+1})$ with $\chi(y)=1$ for $|y|<\frac{1}{2}$ and $\chi(y)=0$ for $|y|>1$.
Then, with the $X_{i}$ acting in $w=(t,x,\xi)=\exp_{v}(y)$, we can write
(3.5) $\Bigl(X_{0}-\sum_{i=1}^{n}X_{i}^{2}\Bigr)T_{K}f(w)=f(w)+\int R_{N}(v,\Theta_{v}(w))f(v)\,dv,$
Here $R_{N}(v,y)$ is a time-independent finite sum of the form
(3.6) $R_{N}(v,y)=\!\!\sum_{\operatorname{order}(\gamma)\geq N-1}r_{\gamma}(v,y)K_{\gamma}(y),\;\;r_{\gamma}(v,y)\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},v}.$
We will establish smoothing estimates for such $R_{N}$ in Section 5.
By hypoellipticity of $X_{0}-\sum_{i=1}^{n}X_{i}^{2}$ and its transpose, a right parametrix is necessarily also a left parametrix (with care taken concerning the order of the remainder), but we can easily modify the above to directly produce a left parametrix.
For each $N$ and $w$ we want to produce a kernel $K(w,\cdot)$ such that
$\int K(w,-y)\Bigl(X_{0}f-\sum_{i=1}^{n}X_{i}^{2}f\Bigr)(\exp_{w}(y))\,\chi(y)\,dy\\ =f(w)+\int R_{N}(w,-y)f(\exp_{w}(y))\,dy.$
If we expand $X_{0}-\sum_{i=1}^{n}X_{i}^{2}$ to order $N$ as above, we seek kernels $K_{m}(y)$ of order m-m as in Theorem 2.5 so that, with $P_{j}^{t}$ the transpose of $P_{j}$ with respect to $dy$,
$\Bigl(P_{0}^{t}(-y,-\partial_{y})+\sum_{j=1}^{N}P_{j}^{t}(-y,-\partial_{y})\Bigr)\Bigl(K_{0}(y)+\sum_{m=1}^{N}K_{m}(y)\Bigr)=\delta(y)+R_{N}(y).$
Observe that
$P_{0}^{t}(-y,-\partial_{y})=\Bigl(\partial_{0}+{\tfrac{1}{2}}\sum_{i=1}^{n}y_{i}\partial_{i+n}\Bigl)-\sum_{i=1}^{n}\Bigl(\partial_{i}-{\tfrac{1}{2}}y_{0}\partial_{i+n}\Bigr)^{2},$
which is the conjugation of $P_{0}(y,\partial_{y})$ under $y^{\prime\prime}\rightarrow-y^{\prime\prime}$. This conjugation preserves both $K_{0}(y)$ and $\delta(y)$, hence $P_{0}^{t}(-y,-\partial_{y})K_{0}(y)=\delta(y)$. This conjugation also preserves the order of each $P_{j}$. Thus if we set
$\tilde{K}_{m}(y)=K_{m}(y_{0},y^{\prime},-y^{\prime\prime}),$
then with $\tilde{P}_{j}^{t}$ the conjugation of $P_{j}^{t}$ under $y^{\prime\prime}\rightarrow-y^{\prime\prime}$, we require that
$\Bigl(P_{0}(y,\partial_{y})+\sum_{j=1}^{N}\tilde{P}_{j}^{t}(-y,-\partial_{y})\Bigr)\Bigl(K_{0}(y)+\sum_{m=1}^{N}\tilde{K}_{m}(y)\Bigr)=\delta(y)+\tilde{R}_{N}(y).$
We can iteratively construct such kernels following Theorem 2.5, noting that $\tilde{K}_{\gamma}(y)=(-1)^{|\alpha^{\prime\prime}|}K_{\gamma}(y)$. We summarize the result in the following.
Theorem 3.5.
For each $m\geq 1$ there are a finite collection of $t$-independent functions $b_{\gamma}(w,y_{0})\in C_{y_{0}}C^{\infty}_{w}(\mathbb{R}^{2n+2})$ where $\operatorname{order}(\gamma)=m$, such that
$K(w,y)=K_{0}(y)+\sum_{m=1}^{N}\sum_{\operatorname{order}(\gamma)=m}b_{\gamma}(w,y_{0})\,K_{\gamma}(y),$
satisfies
$\int K(w,-y)\Bigl(X_{0}f-\sum_{i=1}^{n}X_{i}^{2}f\Bigr)(\exp_{w}(y))\,\chi(y)\,dy\\ =f(w)+\int R_{N}(w,-y)f(\exp_{w}(y))\,dy,$
where $R_{N}(w,y)$ is of the form (3.6).
4. The case $0<\epsilon\leq 1$
In this section we construct parametrices for $X_{0}-\epsilon^{2}\sum_{i=1}^{n}X_{i}^{2}$, $\epsilon\in(0,1]$. Consider the scaled kernels
$K_{\gamma,\epsilon}(y)=\epsilon^{-2n}K_{\gamma}(y_{0},\epsilon^{-1}y^{\prime},\epsilon^{-1}y^{\prime\prime}).$
The scaling is such that $K_{0,\epsilon}(y_{0},\cdot)$ converges to $\delta(y^{\prime},y^{\prime\prime})$ as $y_{0}\rightarrow 0+$, and
$\Bigl(Y_{0}-\epsilon^{2}\sum_{j=1}^{n}Y_{j}^{2}\Bigr)K_{0,\epsilon}(y)=\delta(y).$
More generally,
(4.1) $y^{\alpha}(\partial_{y_{0}},\epsilon\partial_{y^{\prime}},\epsilon\partial_{y^{\prime\prime}})^{\beta}K_{\gamma,\epsilon}(y)=\epsilon^{|\alpha^{\prime}|+|\alpha^{\prime\prime}|}\bigl(y^{\alpha}\partial_{y}^{\beta}K_{\gamma}\bigr)_{\epsilon}(y).$
We note that, with the notation of Lemma 2.3
$\int K_{0,\epsilon}(z^{-1}y)f(z_{0})K_{0,\epsilon}(z)\,dz=F(y_{0})K_{0,\epsilon}(y),$
which can be seen by a change of variables and observing that $\epsilon$-scaling intertwines with group multiplication.
The expansion of $X_{0}-\epsilon^{2}\sum_{j=1}^{n}X_{j}^{2}$ in exponential coordinates $y$ is
$\biggl(Y_{0}-\epsilon^{2}\sum_{i=1}^{n}Y_{i}^{2}\biggr)+\sum_{j\geq 1}P_{j}(\epsilon,y,\partial_{y})$
where $P_{j}$ is a sum of terms with $\operatorname{order}(y^{\alpha}\partial_{y}^{\beta})=2-j$ of the form
(4.2) $P_{j}(\epsilon,y,\partial_{y})=\sum_{|\beta|\leq 2}\sum_{\alpha>0}\epsilon^{2-|\beta|}f_{\alpha,\beta}(y_{0})y^{\alpha}(\epsilon\partial_{y^{\prime}})^{\beta^{\prime}}(\epsilon\partial_{y^{\prime\prime}})^{\beta^{\prime\prime}}$
Here $f_{\alpha,\beta}(y_{0})\in C^{0}_{y_{0}}C^{\infty}_{w}$. The remainder term of order 1N1-N is of the same form with coefficients $f_{\alpha,\beta}(y)\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$.
By (4.1), the iterative procedure of Theorem 2.5 yields the following.
Theorem 4.1.
Suppose given differential operators $P_{j}(\epsilon,y,\partial_{y})$ of the form (4.2). Then there are kernels $K_{m}(\epsilon,y)$, each a finite sum of the form
(4.3) $K_{m}(\epsilon,y)=\sum_{j\geq 0}\sum_{\operatorname{order}(\gamma)=m}b_{j,\gamma}(y_{0})\,\epsilon^{j}K_{\gamma,\epsilon}(y)$
so that
$\Bigl(Y_{0}-\epsilon^{2}\sum_{j=1}^{n}Y_{j}^{2}+\sum_{j=1}^{N}P_{j}(\epsilon,y,\partial_{y})\Bigr)\Bigl(K_{0,\epsilon}(y)+\sum_{m=1}^{N}K_{m}(\epsilon,y)\Bigr)\\ =\delta(y)+R_{N}(\epsilon,y),$
where $R_{N}(\epsilon,y)$ is a finite sum of the form (4.3) with $\operatorname{order}(\gamma)\geq N-1$.
We can use this to give left and right parametrices for $X_{0}-\epsilon^{2}\sum_{i=1}^{n}X_{i}^{2}$. We state the result for the left parametrix, observing that the transpose of an operator $P_{j}$ of the form (4.2) is of the same form.
Theorem 4.2.
For each $m\geq 1$ there are a finite collection of $t$-independent functions $b_{j,\gamma}(w,y_{0})\in C^{0}_{y_{0}}C^{\infty}_{w}$ where $\operatorname{order}(\gamma)=m$, such that
$K(\epsilon,w,y)=K_{0,\epsilon}(y)+\sum_{m=1}^{N}\sum_{\operatorname{order}(\gamma)=m}b_{j,\gamma}(w,y_{0})\,\epsilon^{j}K_{\gamma,\epsilon}(y),$
satisfies
$\int K(\epsilon,w,-y)\Bigl(X_{0}f-\epsilon^{2}\sum_{i=1}^{n}X_{i}^{2}f\Bigr)(\exp_{w}(y))\,\chi(y)\,dy\\ =f(w)+\int R_{N}(\epsilon,w,-y)f(\exp_{w}(y))\,dy,$
where $R_{N}(\epsilon,w,y)$ is a finite sum over $j\geq 0$ and $\operatorname{order}(\gamma)\geq N-1$,
$R_{N}(\epsilon,w,y)=\sum r_{j,\gamma}(w,y)\,\epsilon^{j}K_{\gamma,\epsilon}(y),$
with $r_{j,\gamma}\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$.
5. Operator bounds
We start by observing the following, which holds by a change of variables,
(5.1) $\int\epsilon^{-2n}\bigl|K_{\gamma,\epsilon}(y)\bigr|\,dy^{\prime}\,dy^{\prime\prime}=c_{\gamma}\,y_{0}^{\frac{1}{2}\operatorname{order}(\gamma)}.$
In particular, if $\operatorname{order}(\gamma)>0$ then $\lim_{y_{0}\rightarrow 0}\|K_{\gamma,\epsilon}(y_{0},\cdot)\|_{L^{1}}=0$.
We use this to establish fixed time $L^{p}$ bounds on kernels of nonpositive order, for all $1\leq p\leq\infty$. The left parametrix is an integral kernel $K(\epsilon,w,v)$ where $K$ is a finite sum with $\operatorname{order}(\gamma)\geq 0$, and
$K(\epsilon,w,\exp_{w}(y))=\sum b_{\gamma,\epsilon}(w,y)K_{\gamma,\epsilon}(-y)$
with $b_{\gamma,\epsilon}\in C^{0}_{y_{0}}C^{\infty}_{y^{\prime},y^{\prime\prime},w}$.
Recall from Definition 3.4 that $y=\Theta_{w}(v)$ is the solution to $v=\exp_{w}(y)$, and we use notation $v=(s,z,\zeta)$, $w=(t,x,\xi)$. For fixed $w$, respectively fixed $v$, the maps $y\rightarrow v$ and $y\rightarrow w$ are 1-1 and the Jacobian factors satisfy
$\biggl|\frac{D\Theta_{w}(v)}{Dv}\biggr|=\det(B)^{-2}=\biggl|\frac{D\Theta_{w}(v)}{Dw}\biggr|$
From (5.1), and the fact that $y_{0}=v_{0}-w_{0}=s-t$, we deduce that
Lemma 5.1.
For each $\gamma$ there is $C_{\gamma}$ so that
$\int\bigl|K_{\gamma,\epsilon}(-\Theta_{w}(v))\bigr|\,\delta(v_{0})\,dv$ $\leq C_{\gamma}\,|w_{0}|^{\frac{1}{2}\operatorname{order}(\gamma)},$
$\int\bigl|K_{\gamma,\epsilon}(-\Theta_{w}(v))\bigr|\,\delta(w_{0})\,dw$ $\leq C_{\gamma}\,|v_{0}|^{\frac{1}{2}\operatorname{order}(\gamma)}.$
We next observe that for $\alpha\in\mathbb{Z}_{+}^{2n}$, by (3.2) one can express
$(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}K_{\gamma,\epsilon}(-\Theta_{w}(v))=\sum_{\theta}c_{\theta}K_{\theta,\epsilon}(-\Theta_{w}(v))$
where the finite sum is over $\operatorname{order}(\theta)=\operatorname{order}(\gamma)-\operatorname{order}(\alpha)$.
By the Schur test, together with time translation invariance of $\Theta_{w}(v)$, we obtain the following fixed time mapping properties, where for a function $f\in L^{1}_{\mathit{loc}}(\mathbb{R}^{2n})$ we use the notation
$(\delta\otimes f)(v)=\delta(v_{0})\,f(v^{\prime},v^{\prime\prime}).$
Corollary 5.2.
For all $\gamma$, and all $\alpha\in\mathbb{Z}_{+}^{2n}$, there is $C_{\alpha,\gamma}(T)$ so that, for all $1\leq p\leq\infty$ and all $0<\epsilon\leq 1$, and $t=w_{0}\in[0,T]$,
$T_{\gamma,\epsilon}f(w)=\int K_{\gamma,\epsilon}(-\Theta_{w}(v))\,(\delta\otimes f)(v)\,dv$
satisfies
$\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}(T_{\gamma,\epsilon}f)(t,\cdot)\|_{L^{p}(\mathbb{R}^{2n})}\leq C_{\alpha,\gamma}(T)\,t^{\frac{1}{2}\operatorname{order}(\gamma)-\frac{1}{2}\operatorname{order}(\alpha)}\|f\|_{L^{p}(\mathbb{R}^{2n})}.$
For fixed-time derivative estimates of integer order, we need to work with multiples of order 3, since this is the order of both $\partial_{x}$ and $(\partial_{\xi})^{3}$. The key result (for $\epsilon=1$) is the following.
Lemma 5.3.
Suppose that $P(y,\partial_{y^{\prime}},\partial_{y^{\prime\prime}})$ is a polynomial differential operator of order at most $3m+\operatorname{order}(\gamma)$, where $m\in\mathbb{Z}_{+}$ and $\operatorname{order}(\gamma)\geq 0$. Then one can write
$\bigl(P(y,\partial_{y^{\prime}},\partial_{y^{\prime\prime}})K_{\gamma}\bigr)(-y)=\sum_{\beta,\theta}c_{\beta,\theta}\,(Y^{\beta})^{t}\bigl(K_{\theta}(-y)\bigr),$
where $\operatorname{order}(\beta)\leq 3m$, $\beta_{0}=0$, and $\operatorname{order}(\theta)\geq 0$.
Proof.
Observe that $(Y^{\prime})^{t}f(-y)=(\partial_{y^{\prime}}f-{\tfrac{1}{2}}y_{0}\partial_{y^{\prime\prime}}f)(-y)$, and $(Y^{\prime\prime})^{t}f(-y)=(\partial_{y^{\prime\prime}}f)(-y)$. Writing $\partial_{y^{\prime}}$ and $\partial_{y^{\prime\prime}}$ in terms of $(Y^{\prime})^{t}$ and $(Y^{\prime\prime})^{t}$ expands $P(y,\partial_{y^{\prime}},\partial_{y^{\prime\prime}})K_{\gamma}$ as a sum, with $|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3m+\operatorname{order}(\sigma)$ and $\sigma\geq\gamma$,
$\sum c_{\alpha,\sigma}\,(\partial_{y^{\prime}}-{\tfrac{1}{2}}y_{0}\partial_{y^{\prime\prime}})^{\alpha^{\prime}}(\partial_{y^{\prime\prime}})^{\alpha^{\prime\prime}}K_{\sigma}.$
For a term with $|\alpha^{\prime\prime}|\geq m$, one can factor out $m$ powers of $\partial_{y^{\prime\prime}}$ and let the remaining derivatives fall on $K_{\sigma}$, which yields a sum of kernels of the form $K_{\theta}$ with $\operatorname{order}(\theta)\geq 0$.
A term with $|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3m$ is already in the desired form. For a term with $|\alpha^{\prime\prime}|<m$ and $|\alpha^{\prime}|+3|\alpha^{\prime\prime}|>3m$ we factor out $(\partial_{y^{\prime\prime}})^{\alpha^{\prime\prime}}$ and $3m-3|\alpha^{\prime\prime}|$ powers of $(\partial_{y^{\prime}}-{\tfrac{1}{2}}y_{0}\partial_{y^{\prime\prime}})$; the remaining powers applied to $K_{\sigma}$ leads to kernels $K_{\theta}$ with $\operatorname{order}(\theta)\geq 0$. ∎
By a dilation in $\epsilon$, together with (3.4) for $N$ sufficiently large, we obtain the following.
Corollary 5.4.
Suppose that $P(\epsilon,y,\epsilon\partial_{y^{\prime}},\epsilon\partial_{y^{\prime\prime}})$ is polynomial in each term, with $\operatorname{order}(P)\leq 3m+\operatorname{order}(\gamma)$ where $m\in\mathbb{Z}_{+}$. Then one can write
$\bigl(P(\epsilon,y,\epsilon\partial_{y^{\prime}},\epsilon\partial_{y^{\prime\prime}})K_{\gamma,\epsilon}\bigr)(-y)=\sum_{j,\beta,\theta}c_{j,\beta,\theta}\,\epsilon^{j}(\epsilon^{|\beta|}Y^{\beta})^{t}\bigl(K_{\theta,\epsilon}(-y)\bigr),$
where $j\geq 0$, $\operatorname{order}(\beta)\leq 3m$, $\beta_{0}=0$, and $\operatorname{order}(\theta)\geq 0$.
Furthermore, with $X_{j}$ the representation of $X_{j}$ in exponential coordinates at a given point $w$, one can write
(5.2) $\bigl(P(\epsilon,y,\epsilon\partial_{y^{\prime}},\epsilon\partial_{y^{\prime\prime}})K_{\gamma,\epsilon}\bigr)(-y)=\sum_{j,\beta,\theta}c_{j,\beta,\theta}(w,y_{0})\,\epsilon^{j}(\epsilon^{|\beta|}X^{\beta})^{t}\bigl(K_{\theta,\epsilon}(-y)\bigr)\\ +\sum_{j,\sigma}r_{j,\sigma}(w,y)\epsilon^{j}K_{\sigma,\epsilon}(-y)$
for functions $c_{j,\beta,\gamma}\in C^{0}_{y_{0}}C^{\infty}_{w}$, with the same conditions on $j$, $\beta$ and $\theta$, and $r_{j,\sigma}\in C^{0}_{y_{0}}C^{\infty}_{w,y^{\prime},y^{\prime\prime}}$, with $j\geq 0$ and $\operatorname{order}(\sigma)\geq 0$.
Theorem 5.5.
Suppose that $\operatorname{order}(\gamma)\geq 0$, and let
$T_{\gamma,\epsilon}f(w)=\int K_{\gamma,\epsilon}(-\Theta_{w}(v))\,(\delta\otimes f)(v)\,\chi(\Theta_{w}(v))\,dv$
Then if $|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3m+\operatorname{order}(\gamma)$, $m\in\mathbb{Z}_{+}$, there is $C_{\gamma,m}$ so that, for all $1\leq p\leq\infty$ and all $0<\epsilon\leq 1$,
$\sup_{t\geq 0}\,\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}T_{\gamma,\epsilon}f(t,\cdot)\bigr\|_{L^{p}(\mathbb{R}^{2n})}\\ \leq C_{\gamma,m}\sum_{|\beta^{\prime}|+3|\beta^{\prime\prime}|\leq 3m}\|(\epsilon\partial_{\xi})^{\beta^{\prime}}(\epsilon\partial_{x})^{\beta^{\prime\prime}}f\|_{L^{p}}.$
Proof.
Applying (3.2) we write
$(\epsilon X^{\prime})^{\alpha^{\prime}}(\epsilon X^{\prime\prime})^{\alpha^{\prime\prime}}K_{\gamma,\epsilon}(-\Theta_{w}(v))=\bigl((\epsilon Y^{\prime})^{\alpha^{\prime}}(\epsilon Y^{\prime\prime})^{\alpha^{\prime\prime}}K_{\gamma,\epsilon}\bigr)(-\Theta_{w}(v)).$
We can apply (5.2) since the transpose in $dv$ is the same as in $dy$,
$(\epsilon X^{\prime})^{\alpha^{\prime}}(\epsilon X^{\prime\prime})^{\alpha^{\prime\prime}}T_{\gamma,\epsilon}f(w)\\ =\sum c_{j,\theta,\beta}(w,w_{0})\,\epsilon^{j}\int K_{\theta,\epsilon}(-\Theta_{w}(v))\,(\delta\otimes(\epsilon X)^{\beta}f)(v)\,\chi(\Theta_{w}(v))\,dv\\ +\int R(\epsilon,w,v,-\Theta_{w}(v))\,(\delta\otimes f)(v)\,dv$
where the sum is over $\operatorname{order}(\beta)\leq 3m$, $\beta_{0}=0$, and $\operatorname{order}(\theta)\geq 0$. $R$ is a sum of terms $\epsilon^{j}r_{j,\sigma}(w,\Theta_{w}(v))K_{\sigma,\epsilon}(-\Theta_{w}(v))$ with $\operatorname{order}(\sigma)\geq 0$. The bound of the theorem now follows from Corollary 5.2, since we are restricted to a unit time interval by the support of $\chi$. ∎
6. Proof of Theorem 1.1
For simplicity we consider data $u_{0}(x,\xi)\in\mathcal{S}(\mathbb{R}^{2n})$; the a priori estimates will allow the extension of Theorem 1.1 to spaces of functions with $L^{p}$ bounds on a finite number of derivatatives. We use the following two results.
Take $Q$ as in (1.2), and let $u=e^{tQ}u_{0}$. Then
Maximum Principal: $\sup_{t>0}\sup_{x,\xi}u(t,x,\xi)\leq\sup_{x,\xi}u_{0}(x,\xi).$
Mass Conservation: for $t>0$, $\;\int_{\mathbb{R}^{2n}}u(t,x,\xi)\,dx\,d\xi=\int_{\mathbb{R}^{2n}}u_{0}(x,\xi)\,dx\,d\xi.$
Both of these follow as for the standard heat equation; for completeness we include the proofs here. For the maximum principle, it suffices to work on a strip $t\in[0,T]$ for $T<\infty$. Given $c>0$ consider the function $u_{c}=u-ct$, so $(\partial_{t}-Q)u_{c}=-c<0$. At a local maximum of $u_{c}$ over $0\leq t\leq T$ we must have $\nabla_{x}u=0=\nabla_{\xi}u$, and $X_{j}^{2}u\leq 0$ for $j\geq 1$. Additionally $\partial_{t}u\geq 0$ if $t\neq 0$. This leads to a contradiction if $t\neq 0$. Thus $u_{c}$ attains it maximum at $t=0$. Since this holds for all $c>0$, it follows that $u$ attains its maximum at $t=0$.
Mass conservation follows by writing
$\partial_{t}\int u\,dx\,d\xi=\int Qu\,dx\,d\xi.$
The first order terms in $Q$ are divergence free, and the second order terms have constant coefficients, hence we may integrate by parts, using decay of $u$ in $(x,\xi)$, to see that the right hand side vanishes.
Since the equation is linear and real, the maximum principal shows that
$\|e^{tQ}u_{0}\|_{L^{\infty}(\mathbb{R}^{2n})}\leq\|u_{0}\|_{L^{\infty}(\mathbb{R}^{2n})}.$
The same bound holds for the transpose of $Q$, so by duality this also holds for the $L^{1}$ norm. Alternatively, the $L^{1}$ contractivity follows by positivity together with mass conservation. We conclude by interpolation that
(6.1) $\|e^{tQ}u_{0}\|_{L^{p}(\mathbb{R}^{2n})}\leq\|u_{0}\|_{L^{p}(\mathbb{R}^{2n})}\quad\text{for all}\;\;1\leq p\leq\infty.$
Proof of Theorem 1.1.
Given $N$, let $T_{K}$ be a right inverse for $\partial_{t}-Q$ as in (3.5) such that the remainder is of order 3N-3N, and let $g=T_{K}(\delta\otimes u_{0})$. Since $K(v,y)-K_{0,\epsilon}(y)$ is of negative order, and $K_{0,\epsilon}(y_{0},\cdot)\rightarrow\delta(\cdot)$ as $y_{0}\rightarrow 0+$, by (5.1) we have that $\lim_{t\rightarrow 0+}g(t,\cdot)=u_{0}$ in the $L^{p}(\mathbb{R}^{2n})$ norm.
Recall that $X^{\prime}=B^{T}\partial_{\xi}$ and $X^{\prime\prime}=B^{T}\partial_{x}$ with $\det(B)\neq 0$. The coefficients of $K$ are continuous in $y_{0}$ and smooth in $v^{\prime},v^{\prime\prime}$. By Theorem 5.5 we deduce that, for every $m\in\mathbb{Z}_{+}$,
(6.2) $\sup_{t>0}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3m}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}g(t,\cdot)\bigr\|_{L^{p}}\\ \leq C_{N}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3m}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}u_{0}\bigr\|_{L^{p}}.$
Additionally, $(\partial_{t}-Q)g=R_{3N+1}(\delta\otimes u_{0})$ satisfies
(6.3) $\sup_{t>0}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3N}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}(\partial_{t}-Q)g(t,\cdot)\bigr\|_{L^{p}}\leq C_{N}\|u_{0}\|_{L^{p}}.$
Let $h=e^{tQ}u_{0}-g$, so $(\partial_{t}-Q)h=-(\partial_{t}-Q)g$ and $h=0$ for $t\leq 0$. To control $h$, let $K$ be the left parametrix constructed in Theorem 4.2 for $\partial_{t}-Q$ with remainder $R_{3N+1}$ of order 3N-3N, and write
$h(w)=-\int K(\epsilon,w,-y)\bigl(\partial_{t}g-Qg\bigr)(\exp_{w}(y))\,\chi(y)\,dy\\ +\int R_{3N+1}(\epsilon,w,-y)h(\exp_{w}(y))\,dy.$
By (6.1) and (6.2) with $m=0$ we have $\|h(t,\cdot)\|_{L^{p}}\leq C\|u_{0}\|_{L^{p}}$ for $t\geq 0$. If $|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq N$, and $w=(t,x,\xi)$, Theorem 5.5 with $m=0$ then yields
$\sup_{t>0}\,\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}\int R_{3N+1}(\epsilon,w,-y)h(\exp_{w}(y))\,dy\,\bigr\|_{L^{p}(dx\,d\xi)}\\ \leq C\|u_{0}\|_{L^{p}},$
where we use that the kernel of $R_{3N+1}$ is supported in $|y|<1$ so uniform in time $L^{p}$ bounds on $h$ control the integral over $y$. To handle the first term in $h$ we use (6.3) and Theorem 5.5 with $m=3N$. Together we conclude that
$\sup_{t>0}\sum_{|\alpha^{\prime}|+3|\alpha^{\prime\prime}|\leq 3N}\bigl\|(\epsilon\partial_{\xi})^{\alpha^{\prime}}(\epsilon\partial_{x})^{\alpha^{\prime\prime}}h(t,\cdot)\bigr\|_{L^{p}(dx\,d\xi)}\leq C\|u_{0}\|_{L^{p}}.$
Together with (6.2) this establishes the bounds of (1.5) on $u$.
To establish (1.6) we follow the same steps, and note that the derivatives up to order $3N$ of $h$ satisfy uniform $L^{p}$ bounds for $t\geq 0$. Additionally, $g$ is supported in $t\leq 1$ and satisfies the bounds of (1.6) by Corollary 5.2, noting that Lemma 5.1 is invariant under transpose. ∎
7. Applications to classical-quantum correspondence
We show here that certain results from [3] and [7] hold in the subelliptic setting considered in this paper. We consider real valued $V$ satisfying (1.1), and complex valued $\ell_{j}$ satisfying (1.3) with non-degenerate $B$, and define $Q$ by (1.2). Our first application builds upon results of [7] concerning closeness in the trace norm of the quantum and classical evolution of certain semi-classical observables. The symbol class $S^{L^{q}}_{\rho}$ denotes symbols $a(h,x,\xi)$ such that, for all multi-indices $\alpha$, there is $C_{\alpha}$ so that
$\|\partial_{x,\xi}^{\alpha}a\|_{L^{q}}\leq C_{\alpha}h^{\frac{n}{q}-\rho|\alpha|}.$
This space of symbols is topologized by the seminorms
$\|a\|_{N,S^{L^{q}}_{\rho}}=\sum_{|\alpha|\leq N}h^{\rho|\alpha|-\frac{n}{q}}\|\partial_{x,\xi}^{\alpha}a\|_{L^{q}}.$
For $q=1$, as noted in Lemma 2.1 of [7] the connection to the trace norm is given by Theorem 9.3 of [1] and (C.3.1) of [10],
(7.1) $\|\mathrm{Op}_{h}^{w}(a(t))\|_{\mathcal{L}^{1}}\leq C_{n}\|a\|_{2n+1,S^{L^{1}}_{1/2}}$
Theorem 7.1.
Assume that $a\in S^{L^{1}}_{1/2}$, and let $a(t)=e^{tQ}a$. If $A(t)$ satisfies $\partial_{t}A(t)=\mathcal{L}A(t)$ with $A(0)=\mathrm{Op}_{h}^{w}(a)$, there is $C$ depending on $\|a_{0}\|_{N,S^{L^{1}}_{1/2}}$ for some NN, such that
$\|A(t)-\mathrm{Op}_{h}^{w}(a(t))\|_{\mathcal{L}^{1}}\leq Cth^{\frac{1}{2}},\quad 0\leq t<\infty.$
Proof.
The proof follows that of Theorem 2 of [7]. We include the outline of the proof for convenience. Using $A(0)=\mathrm{Op}_{h}^{w}(a(0))$ one writes
$A(t)-\mathrm{Op}_{h}^{w}(a(t))$ $=\int_{0}^{t}e^{(t-s)\mathcal{L}}(\mathcal{L}-\partial_{s})\mathrm{Op}_{h}^{w}(a(s))\,ds$
$=\int_{0}^{t}e^{(t-s)\mathcal{L}}\Bigl(\mathcal{L}\mathrm{Op}_{h}^{w}(a(s))-\mathrm{Op}_{h}^{w}(Qa(s))\Bigr)\,ds.$
Since $e^{t\mathcal{L}}$ is a contraction on the space of trace class operators for $t\geq 0$, it suffices to verify that, uniformly for all $s\in[0,\infty)$,
$\Bigl\|\mathcal{L}\mathrm{Op}_{h}^{w}(a(s))-\mathrm{Op}_{h}^{w}(Qa(s))\Bigr\|_{\mathcal{L}^{1}}\leq Ch^{\frac{1}{2}}.$
Since $\gamma>0$ is fixed, we have $\epsilon=\sqrt{\gamma h/2}\approx h^{\frac{1}{2}}$. Theorem 1.1 thus implies
$\sup_{s\geq 0}\|a(s)\|_{N,S_{1/2}^{L^{1}}}\leq C_{N}\,\|a\|_{N,S_{1/2}^{L^{1}}},\quad\forall N\geq 0.$
The jump operators $L_{j}$ are linear functions of $x$, so the composition rules for the semi-classical Weyl calculus imply that
(7.2) $\mathcal{L}\mathrm{Op}_{h}^{w}(a(s))-\mathrm{Op}_{h}^{w}(Qa(s))=\frac{i}{h}[P,\mathrm{Op}_{h}^{w}(a(s))]-\mathrm{Op}_{h}^{w}(H_{p}a(s)).$
The proof is thus concluded by applying composition results from [10] as in Lemma 2.3 of [7] to obtain, for NN depending on $n$,
$\Bigl\|\frac{i}{h}[P,\mathrm{Op}_{h}^{w}(a(s))]-\mathrm{Op}_{h}^{w}(H_{p}a(s))\Bigr\|_{2n+1,S_{1/2}^{L^{1}}}\leq Ch^{\frac{1}{2}}\|a(s)\|_{N,S_{1/2}^{L^{1}}},$
together with (7.1). ∎
We now turn to results of [3], which concern bounds on the Hilbert-Schmidt norm of the difference, and relate to our estimates for $q=2$. The connection of $S_{\rho}^{L^{2}}$ to the Hilbert-Schmidt norm is given by
(7.3) $\|\mathrm{Op}_{h}^{w}(a(t))\|_{\mathcal{L}^{2}}=(2\pi h)^{-\frac{n}{2}}\|a(t)\|_{L^{2}}=(2\pi)^{-\frac{n}{2}}\|a(t)\|_{0,S^{L^{2}}_{\rho}}.$
In the following theorem the coupling constant $\gamma$ is allowed to depend on $h$, subject to lower bounds of $h^{\frac{1}{3}}$.
Theorem 7.2.
Assume that $\frac{1}{2}\leq\rho\leq\frac{2}{3}$, $a\in S^{L^{2}}_{\rho}$, and let $a(t)=e^{tQ}a$. Assume also that $\gamma=\gamma(h)$ satisfies $h^{2\rho-1}\leq\gamma\leq 1$. Then if $A(t)$ satisfies $\partial_{t}A(t)=\mathcal{L}A(t)$ with $A(0)=\mathrm{Op}_{h}^{w}(a)$, there is $C$ depending on $\|a_{0}\|_{N,S^{L^{2}}_{\rho}}$ for some NN, such that
$\|A(t)-\mathrm{Op}_{h}^{w}(a(t))\|_{\mathcal{L}^{2}}\leq Cth^{2-3\rho},\quad 0\leq t<\infty.$
Proof.
The proof is similar to that of Theorem 7.1. The semigroup $e^{t\mathcal{L}}$ is a contraction on Hilbert-Schmidt operators for $t\geq 0$ by Proposition 4.6 of [3], as $M=0$ in our case, so it suffices to verify that, uniformly for all $s\in[0,\infty)$,
$\Bigl\|\mathcal{L}\mathrm{Op}_{h}^{w}(a(s))-\mathrm{Op}_{h}^{w}(Qa(s))\Bigr\|_{\mathcal{L}^{2}}\leq Ch^{2-3\rho}.$
Since $1\geq\gamma\geq h^{2\rho-1}$ we have $h^{\frac{1}{2}}\geq\epsilon\geq h^{\rho}$. Theorem 1.1 thus implies
$\sup_{s\geq 0}\|a(s)\|_{N,S_{\rho}^{L^{2}}}\leq C_{N}\,\|a\|_{N,S_{\rho}^{L^{2}}},\quad\forall N\geq 0.$
By (7.2) and (7.3) the theorem follows from showing that
$\Bigl\|\frac{i}{h}[P,\mathrm{Op}_{h}^{w}(a(s))]-\mathrm{Op}_{h}^{w}(H_{p}a(s))\Bigr\|_{0,S_{\rho}^{L^{2}}}\leq Ch^{2-3\rho}\|a(s)\|_{N,S_{\rho}^{L^{2}}}$
for some NN depending on $n$, which holds by Lemma 2.2 of [3]. ∎
Acknowledgements
I would like to thank Maciej Zworski for bringing the topic of this paper to my attention, and for his encouragement in pursuing the results obtained herein. I would also like to thank Kevin Li for helpful discussions on the details of his paper [7].